OpenStudy (anonymous):

how do prove this identity: sin^4x -cos^4x=1-2cos^2x

OpenStudy (anonymous):

Do you still need help?

OpenStudy (anonymous):

RS: (sin^2x+cos^2x)-2cos^2x sin^2x-cos^2x Thats how much i can simplify right side..

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

alright.

OpenStudy (anonymous):

So here is how I proved it:

OpenStudy (anonymous):

So, change sinx^4 to (1-cosx^2)^2

OpenStudy (anonymous):

You can expand that and you get 1-2cosx^2+cosx^4-cosx^4=1-2cosx^2

OpenStudy (anonymous):

Do you follow so far?

OpenStudy (anonymous):

@sabika13 Should I continue?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

When you simplify 1-2cosx^2+cosx^4-cosx^4=1-2cosx^2, the left side becomes 1-2cosx^2=1-2cosx^2

OpenStudy (anonymous):

That proves the identity.

OpenStudy (anonymous):

Normally, when you prove an identity, you work on one side. In this problem I worked on the left side.

OpenStudy (anonymous):

wait how does (1-2cosx^2)^2=1-2cos^2+cosx^4?

OpenStudy (anonymous):

To summarize it: You have sin(x)^4-cos(x)^4=1-2cos(x)^2 change sin(x)^4 --->(1-cos(x)^2)^2 You have: (1-cos(x)^2)^2-cos(x)^4=1-2cos(x)^2 Then expand (1-cos(x)^2)^2 on the left side. You get 1-2cos(x)^2+cos(x)^4-cos(x)^4=1-2cos(x)^2

OpenStudy (anonymous):

Any better?

OpenStudy (anonymous):

Then the: cos(x)^4-cos(x)^4 cancel out.

OpenStudy (anonymous):

Hope that helped.

OpenStudy (anonymous):

im having problems expanding (1-cos^2x)^2 shouldnt it be: 1-cos^4x :S

OpenStudy (anonymous):

So, (1-cos^2x)(1-cos^2x). Just foil out.

OpenStudy (anonymous):

ohhh okay, but why is it not: (1-cos^2x)(1+cos^2x)

OpenStudy (anonymous):

It's squared. It's not a difference of squares.

OpenStudy (anonymous):

ohh i think i get it...

OpenStudy (anonymous):

Any number, or equation squared is the same as it times by itself. Example (6)^2= (6)*(6) or (3x-2)^2= (3x-2)(3x-2). Do you get the point?

OpenStudy (anonymous):

You don't change the signs inside the equation.

OpenStudy (anonymous):

yeah so whats difference of squares then (if you dont mind me asking)

OpenStudy (anonymous):

example x^2-4. When you factor it, you get (x-2)(x+2). It's a^2-b^2

OpenStudy (anonymous):

ohhh thankkyou!!you helped me fix a very big misconception inmy head ;P

OpenStudy (anonymous):

yep, no prob.